pls help, I am begging

Where A is autonomous consumption, A = 900 - P. And the planned Aggregate Expenditure function is AEp = 900 - 5P + .8(Yd) + 200. 1. Lets consider the same Keynesian economy we discussed last week. That is, let A = 900 – 5P, MPC = 0.8 and Ip = 200. Price
Answer #1

(1) Ginen, AEp = 900-50 +0.8 YO+ Ip T=0 (since there is no government) we know, Yo = YT - Y = 4-0 (a) Ip= 200 At Equilibrium,[Ip = 200] [Yo=y] > y = 900- 5p + 0.8 Yo + Ip => y = 900 - 5p +0.890 + 200 - y = goo-5p +0.87+200 = 0.2y = 1100-5p => y = 110I inuestment falls Hence by $100. Autonomous